# Planar Turán numbers of three configurations

### v1

@@ Abstract
The planar Tuán number of $H$, denoted by $ex_{\mathcal{P}}(n,H)$, is defined as the maximum number of edges in an $n$-vertex $H$-free planar graph. The exact value of $ex_{\mathcal{P}}(n,H)$ remains a mystery when $H$ is large (for example, $H$ is a long path or a long cycle), while tight bounds have been established for many small planar graphs such as cycles, paths, $\Theta$-graphs and other small graphs formed by a union of them. One representative graph among such union graphs is $K_{1}+L$ where $L$ is a linear forest without isolated vertices. Previous works solved the cases when $L$ is a path or a matching. In this work, we first investigate the planar Turán number of the graph $K_{1}+L$ when $L$ is the disjoint union of a $P_{2}$ and $P_{3}$. Equivalently, $K_{1}+L$ represents a specific configuration formed by combining a $C_{3}$ and a $\Theta_{4}$. We further consider the planar Turán numbers of the all graphs obtained by combining $C_{3}$ and $\Theta_{4}$. Among the six possible such configurations, three have been resolved in earlier works. For the remaining three configurations (including $K_{1}+(P_{2}\dot{\cup}P_{3})$), we derive tight bounds. Furthermore, we completely characterize all extremal graphs for the remaining two of these three cases. Additionally, for two other unsolved cases involving $K_{1}+L$, we establish improved bounds.
@@ 1 Introduction
Given the aforementioned challenges, researchers have shifted their focus toward determining the exact planar Turán numbers of small planar graphs. So far, tight bounds are known only for a few small planar graphs, including short cycles of length up to eight [1, 5, 6], paths of order at most eleven [8], and small $\Theta$-graphs [10] (we use $\Theta_{k}$ to denote a set of graphs that are obtained from a cycle by adding an additional edge joining two non-consecutive vertices. It is clear that $\Theta_{k}$ is a single graph if $k=4,5$). Other studies have examined structures composed of combined cycles, such as unions of triangles sharing vertices or edges [3, 11] and disjoint union of graphs [2, 11]. Among these configurations, one of the most interesting is $K_{1}+L_{t}$, where $L_{t}$ is a linear forest of $t$ vertices (the graph $K_{1}+L_{t}$ is obtained by joining a new vertex to all the vertices of $L_{t}$). Lan, Shi, and Song [9] proved that
Theorem 1.1 ([9]).
For an integer $4\leq t\leq 6$, let $H$ be a graph on $t$ vertices consisting of disjoint paths. Then $ex_{\mathcal{P}}(n,K_{1}+H)\leq\frac{13(t-1)n}{4t-2}-\frac{12(t-1)}{2t-1}$ for all $n\geq t+1$.
It is worth noting that not all the upper bounds obtained are tight. They further determined a tight upper bound for $ex_{\mathcal{P}}(n,K_{1}+2P_{2})$ when $n\geq 5$, and provided an improved upper bound for $ex_{\mathcal{P}}(n,K_{1}+3P_{2})$. Subsequently, Fang, Wang, and Zhai [3] established the tight bound for $ex_{\mathcal{P}}(n,K_{1}+3P_{2})$ and the tight bound of $ex_{\mathcal{P}}(n,K_{1}+P_{t})$ for $3\leq k\leq 6$. In this paper, we extend this line of research by studying graphs of the form $K_{1}+H$. Specifically, we establish a tight bound for $ex_{\mathcal{P}}(n,K_{1}+H)$ when $H=P_{2}\dot{\cup}P_{3}$ is a disjoint union of $P_{2}$ and $P_{3}$.
@@
Dowden [1] determined that $ex_{\mathcal{P}}(n,H_{1})=3n-6$ when $n\geq 6$. Fang, Wang, and Zhai [3] showed that $ex_{\mathcal{P}}(n,H_{2})\leq\frac{8}{3}(n-2)$ with a sharp upper bound. Because $H_{1}$ is a subgraph of $H_{3}$, $ex_{\mathcal{P}}(n,H_{3})=3n-6$ follows directly. In addition to $H_{4}$, we also focus on determining the planar Turán numbers of the remaining two graphs $H_{5}$, and $H_{6}$.
@@
We determine the planar Turán number of $H_{6}$, which is the disjoint union of $C_{3}$ and $\Theta_{4}$, and characterize all extremal graphs. For convenience, we use $M_{t}$ to denote the union of $\left\lfloor\frac{t}{2}\right\rfloor$ pairwise vertex-disjoint edges and $\left\lceil\frac{t}{2}\right\rceil-\left\lfloor\frac{t}{2}\right\rfloor$ isolated vertices. To describe the extremal $H_{6}$-free planar graphs, we need the outerplanar Turán number of $C_{3}$ (denoted $ex_{\mathcal{OP}}(n,C_{3})$), which is the maximum number of edges in an $n$-vertex $C_{3}$-free outerplanar graph. Fang and Zhai [4] proved that $ex_{\mathcal{OP}}(n,C_{3})=\frac{3n-4}{2}$ for each even integer $n\geq 4$.
@@
If $n\geq 174$, then $ex_{\mathcal{P}}(n,C_{3}\dot{\cup}\Theta_{4})=\left\lfloor\frac{5n}{2}\right\rfloor-4$. Moreover, if $n$ is even, then the extremal planar graph is a copy of $M_{n-2}+K_{2}$; if $n$ is odd, then the extremal planar graph is either a copy of $M_{n-2}+K_{2}$ or $O+\{u\}$, where $O$ represents a $C_{3}$-free outerplanar graph of even order with $ex_{\mathcal{OP}}(n-1,C_{3})=\lfloor\frac{3n}{2}\rfloor-3$ edges.
Let $L^{\prime}_{k}$ denote a linear forest of order $k$ without isolated vertices. Authors in [1, 3, 9] determined $ex_{\mathcal{P}}(n,K_{1}+L^{\prime}_{k})$ for all $L^{\prime}_{k}$ except for the two cases $L^{\prime}_{6}=P_{2}\dot{\cup}P_{4}$ and $L^{\prime}_{6}=2P_{3}$. We give a lower bound and an upper bound for $ex_{\mathcal{P}}(n,K_{1}+H)$ when $H\in\{P_{2}\cup P_{4},2P_{3}\}$, which offers a slight improvement to the upper bounds provided by Lan, Shi and Song (Theorem 1.1) for these scenarios.
Theorem 1.5.
Let $H\in\{P_{2}\cup P_{4},2P_{3}\}$, then $ex_{\mathcal{P}}(n,K_{1}+H)\leq\frac{41n}{14}-\frac{36}{7}$. In addition, if $12|n$, then $ex_{\mathcal{P}}(n,K_{1}+H)\geq\frac{11n}{4}-3$.
The rest of this paper is organized as follows. In Section 2, we introduce some concepts and notation on planar graphs. In Sections 3, 4 and 5, we establish sharp bounds on the planar Turán number of $H_{4},H_{5}$ and $H_{6}$, respectively, and characterize all extremal planar graphs for the latter two graphs. In Section 6, we give a short proof of Theorem 1.5.
@@ 2 Preliminaries
For each $H$-free TB, we can transform it into a new TB by regarding all holes whose facial cycles are $C_{3}$s with corresponding 3-faces, such that the new TB maintains $H$-free. We call such a new TB a solid TB (i.e., a solid TB contains no holes with facial cycle $C_{3}$). In the proofs of Theorems 1.2 and 1.3, we will estimate the maximum triangle-density among all $H$-free TBs and TCs (where $H$ is either $H_{4}$ or $H_{5}$). For this purpose, it suffices to estimate the maximum triangle-density among all $H$-free solid TBs and all $H$-free TCs composed of solid TBs. See Figure 2 as an example, the left side is a plane graph, whose TBs are $B_{1},B_{2}$ and $B_{3}$ (see the Figure 2 (2), two gray faces are holes). $B_{1}^{\prime}$ is a solid TB obtained from $B_{1}$ (see the Figure 2 (3)).
@@
If $\Delta_{I}(G_{E})\leq 9$, then by Ineq. (8), $e(G_{E})=\lfloor\frac{5n}{2}\rfloor-4$ if and only if $|E_{I}(G_{E})|=\frac{n}{2}+4$. By Lemma 5.3, we conclude that $e(G_{E})=\lfloor\frac{5n}{2}\rfloor-4$ implies $G_{E}$ contains a spanning subgraph $G^{\prime}$ that is a copy of $\left(\left\lfloor\frac{n-2}{2}K_{2}\right\rfloor\right)+K_{2}$. Without loss of generality, assume that $G^{\prime}=xy+M$, where $M$ is a matching of size $\lfloor(n-2)/2\rfloor$. If $n$ is even, then $e(G^{\prime})=\lfloor\frac{5n}{2}\rfloor-4=e(G)$, and hence $G=G^{\prime}$ is a copy of $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$. If $n$ is odd, then $|V(G)-V(G^{\prime})|=1$ and $|E(G)-E(G^{\prime})|=2$ (say $V(G)-V(G^{\prime})=\{z\}$ and $E(G)-E(G^{\prime})=\{e_{1},e_{2}\}$). It is clear neither $e_{1}$ nor $e_{2}$ belongs to $G[V(G^{\prime})]$, for otherwise there is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction. Hence, we can assume that $e_{1}=z_{1}z$ and $e_{2}=z_{2}z$. Then $\{z_{1},z_{2}\}=\{x,y\}$, for otherwise there is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction. Therefore, $G=G^{\prime}\cup\{z_{1}z,z_{2}z\}$ is a copy of $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$. On the other hand, $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$ is $\Theta_{4}\dot{\cup}C_{3}$-free obviously.
@@
6 Proof of Theorem 1.5
Let $\mathcal{G}$ be the family of maximal $K_{1}+H$-free plane graphs, where $H\in\{P_{2}\dot{\cup}P_{4},2P_{3}\}$. We first present several preliminary lemmas.
Lemma 6.1.
There is a $G\in\mathcal{G}$ such that $G$ is 2-connected.
Proof.
We claim that $G$ is connected; otherwise, adding a bridge would create a $(K_{1}+H)$-free planar graph with more edges, a contradiction. Assume, to the contrary, that every graph in $\mathcal{G}$ contains cut-vertices. We chose an extremal graph $G$ in $\mathcal{G}$ such that the maximum ordered block $D^{*}$ is as large as possible. Since $G$ is not $2$-connected, a vertex $v\in V(D^{*})$ is a cut-vertex of $G$.
It is clear that $D^{*}$ is $2$-connected; otherwise $G$ would be a tree, contradicting the maximality of $e(G)$. Hence, the boundary of $D^{*}$ is a cycle, denoted by $C^{*}$. Let $u$ be a neighbor of $v$ in $C^{*}$. It is clear that the block-tree of $D^{*}-\{uv\}$ a path such that
  1. 1.
    $u,v$ belong to the two leaf blocks, respectively.
  2. 2.
    neither $u$ nor $v$ is a cut-vertex of $D^{*}-\{uv\}$.
Because $v$ is a cut-vertex of $G$, there exists a block $D\neq D^{*}$ in $G$. Choose a vertex $w\neq v$ on the boundary of $D$, and let $P$ be a $vw$-path in $D$. Construct $G^{\prime}=G-\{uv\}+\{uw\}$. The edge $uv$ is not in any triangle in $G^{\prime}$, and since every edge in $K_{1}+H$ is in a triangle, $G^{\prime}$ remains $K_{1}+H$-free. Since $e(G)=e(G^{\prime})$, $G^{\prime}$ is also an extremal plane graph in $\mathcal{G}$. However, $R=(D^{*}-\{uv\})\cup P\cup\{uw\}$ is 2-connected in $G^{\prime}$ and $|R|>|D^{*}|$. Thus, $G^{\prime}$ contains a block strictly larger than $D^{*}$, contradicting the maximality of $D^{*}$. Consequently, $G$ is $2$-connected. ∎
Lemma 6.2.
Let $G$ be a 2-connected plane graph that is either $(K_{1}+(P_{2}\cup P_{4}))$-free or $(K_{1}+2P_{3})$-free. Then for any vertex $v$ with $d(v)\geq 6$ in $G$, $v$ is adjacent to at least $\frac{d(v)}{3}-1/3$ non-3-faces.
Proof.
Let $v$ be a vertex with $d(v)\geq 6$. We consider two cases based on the forbidden subgraph condition:
Case 1: $G$ is $(K_{1}+(P_{2}\dot{\cup}P_{4}))$-free.
Since $G$ is $(K_{1}+(P_{2}\dot{\cup}P_{4}))$-free, each induced path in $G[N(v)]$ has order at most 5. If $G[N(v)]$ contains a path $P$ with $4\leq|P|\leq 5$, then $G[N(v)\setminus V(P)]$ must be a star. This ensures that $v$ is incident to at least $d(v)-|P|+1\geq\frac{d(v)}{3}$ non-3-faces for $d(v)\geq 6$. If every induced path in $G[N(v)]$ has order at most 3, then $v$ is trivially incident to at least $\frac{d(v)}{3}$ non-3-faces.
Case 2: $G$ is $(K_{1}+2P_{3})$-free.
For this case, each path in $G[N(v)]$ has order at most 5. If $G[N(v)]$ contains a path $P$ with $3\leq|P|\leq 5$, then the remaining graph $G[N(v)\setminus E(P)]$ can only contain paths of order at most 2. This structure guarantees that $v$ is incident to at least $\lceil\frac{d(v)-|P|}{2}\rceil+1\geq\frac{d(v)}{3}-1/3$ non-3-faces for $d(v)\geq 6$. If all induced paths in $G[N(v)]$ have order at most 2, then $v$ is incident to at least $\lceil\frac{d(v)}{2}\rceil\geq\frac{d(v)}{3}$ non-3-faces. ∎
Proof of Theorem 1.5 For the lower bound, we construct a $(K_{1}+H)$-free plane graph on $n$ vertices with $\frac{11n}{4}-3$ edges. Let $n=12k$ for an integer $k\geq 1$. We define the base graph $Q_{1}$ as shown in Figure 11(a). For $k\geq 2$, the graph $Q_{k}$ is constructed recursively by embedding $Q_{k-1}$ into the central triangle of the structure in Figure 11(b), such that the three vertices on the center triangle and the three vertices on the outer boundary of $Q_{k-1}$ are connected by three edges that forms a matching. It is straightforward to verify that each $Q_{k}$ is $(K_{1}+H)$-free and contains exactly $\frac{11n}{4}-3$ edges.
(a) $Q_{1}$
$F_{k-1}$
(b) $Q_{k}$
Figure 11: The construction of $F_{k}$.
We now establish the upper bound. By Lemma 6.1, we may assume $G\in\mathcal{G}$ is 2-connected. Applying a discharging argument to bound the number of missing chords in $D$, we initially assign each $k$-face a charge of k3k-3. This charge is then equally distributed to its incident vertices, yielding at least $\frac{k-3}{k}\geq\frac{1}{4}$ per vertex. Lemma 6.2 further guarantees that each vertex $v$ with $d(v)\geq 6$ is incident to at least $\frac{d(v)}{3}-1/3$ non-3-faces, enabling it to receive a total charge of at least $\frac{1}{4}(\frac{d(v)}{3}-1/3)$. Consequently, we derive
$3n-6-e(G)\geq\frac{1}{4}\sum_{\{v:d(v)\geq 6\}}\left\lfloor\frac{d(v)}{3}\right\rfloor\geq\frac{1}{4}\sum_{\{v:d(v)\geq 6\}}\left(\frac{d(v)}{3}-1/3\right).$
Furthermore, since $5x+\sum_{\{v:d(v)\geq 6\}}d(v)\geq 2e(G)$, where $x$ denotes the number of vertices of degree at most 5, we obtain:
$2e-5x\leq\sum_{\{v:d(v)\geq 6\}}d(v)\leq 36n-72-12e(G)+(n-x).$
Hence, $e(G)\leq\frac{37n-72+4x}{14}\leq\frac{41n}{14}-\frac{36}{7}$, as desired. $\Box$
Based on the construction for the lower bound in Theorem 1.5, a natural question arises:
Question. Is the lower bound in Theorem 1.5 tight when $14|n$?

### v2

@@ Abstract
The planar Tuán number of $H$, denoted by $ex_{\mathcal{P}}(n,H)$, is defined as the maximum number of edges in an $n$-vertex $H$-free planar graph. The exact value of $ex_{\mathcal{P}}(n,H)$ remains a mystery when $H$ is large (for example, $H$ is a long path or a long cycle), while tight bounds have been established for many small planar graphs such as cycles, paths, $\Theta$-graphs and other small graphs formed by a union of them. One representative graph among such union graphs is $K_{1}+L$ where $L$ is a linear forest without isolated vertices. Previous works solved the cases in which $L$ is a path or a matching, or satisfies $|L|\geq 7$. In this work, we first investigate the planar Turán number of the graph $K_{1}+L$ when $L$ is the disjoint union of a $P_{2}$ and $P_{3}$. Equivalently, $K_{1}+L$ represents a specific configuration formed by combining a $C_{3}$ and a $\Theta_{4}$. We further consider the planar Turán numbers of the all graphs obtained by combining $C_{3}$ and $\Theta_{4}$. Among the six possible such configurations, three have been resolved in earlier works. For the remaining three configurations (including $K_{1}+(P_{2}\dot{\cup}P_{3})$), we derive tight bounds. Furthermore, we completely characterize all extremal graphs for the remaining two of these three cases.
@@ 1 Introduction
Given the aforementioned challenges, researchers have shifted their focus toward determining the exact planar Turán numbers of small planar graphs. So far, tight bounds are known only for a few small planar graphs, including short cycles of length up to seven [1, 5, 6], paths of order at most eleven [8], and small $\Theta$-graphs [10] (we use $\Theta_{k}$ to denote a set of graphs that are obtained from a cycle by adding an additional edge joining two non-consecutive vertices. It is clear that $\Theta_{k}$ is a single graph if $k=4,5$). Other studies have examined structures composed of combined cycles, such as unions of triangles sharing vertices or edges [3, 11] and disjoint union of graphs [2, 11]. Among these configurations, one of the most interesting is $K_{1}+L_{t}$, where $L_{t}$ is a linear forest of $t$ vertices (the graph $K_{1}+L_{t}$ is obtained by joining a new vertex to all the vertices of $L_{t}$). Lan, Shi, and Song [9] proved that ffor an integer $4\leq t\leq 6$, let $H$ be a graph on $t$ vertices consisting of disjoint paths. Then $ex_{\mathcal{P}}(n,K_{1}+H)\leq\frac{13(t-1)n}{4t-2}-\frac{12(t-1)}{2t-1}$ for all $n\geq t+1$. It is worth noting that not all the upper bounds obtained are tight. They further determined a tight upper bound for $ex_{\mathcal{P}}(n,K_{1}+2P_{2})$ when $n\geq 5$, and provided an improved upper bound for $ex_{\mathcal{P}}(n,K_{1}+3P_{2})$. Subsequently, Fang, Wang, and Zhai [3] established the tight bound for $ex_{\mathcal{P}}(n,K_{1}+3P_{2})$ and the tight bound of $ex_{\mathcal{P}}(n,K_{1}+P_{t})$ for $3\leq k\leq 6$. In this paper, we extend this line of research by studying graphs of the form $K_{1}+H$. Specifically, we establish a tight bound for $ex_{\mathcal{P}}(n,K_{1}+H)$ when $H=P_{2}\dot{\cup}P_{3}$ is a disjoint union of $P_{2}$ and $P_{3}$.
@@
Dowden [1] determined that $ex_{\mathcal{P}}(n,H_{1})=3n-6$ when $n\geq 6$. Fang, Wang, and Zhai [3] showed that $ex_{\mathcal{P}}(n,H_{2})\leq\frac{8}{3}(n-2)$ with a sharp upper bound. Because $H_{1}$ is a subgraph of $H_{3}$, $ex_{\mathcal{P}}(n,H_{3})=3n-6$ follows directly. In addition to $H_{4}$, we also focus on determining the planar Turán numbers of the remaining two graphs $H_{5}$ and $H_{6}$.
@@
We determine the planar Turán number of $H_{6}$, which is the disjoint union of $C_{3}$ and $\Theta_{4}$, and characterize all extremal graphs. For this purpose, we first introduce two graphs. We use $M_{t}$ to denote the union of $\left\lfloor\frac{t}{2}\right\rfloor$ pairwise vertex-disjoint edges and $\left\lceil\frac{t}{2}\right\rceil-\left\lfloor\frac{t}{2}\right\rfloor$ isolated vertices. For odd $n$, let $K_{2}\vee M_{n-2}$ denote the graph obtained from $K_{2}+M_{n-3}$ by adding an additional vertex $u$ and two edges $uv_{1},uv_{2}$, where $v_{1}$ and $v_{2}$ are endpoints of two arbitrary edges in $M_{n-2}$, respectively. To describe the extremal $H_{6}$-free planar graphs, we also need the notion outerplanar Turán number of $C_{3}$ (denoted $ex_{\mathcal{OP}}(n,C_{3})$), which is the maximum number of edges in an $n$-vertex $C_{3}$-free outerplanar graph. Fang and Zhai [4] proved that $ex_{\mathcal{OP}}(n,C_{3})=\frac{3n-4}{2}$ for each even integer $n\geq 4$.
@@
If $n\geq 174$, then $ex_{\mathcal{P}}(n,C_{3}\dot{\cup}\Theta_{4})=\left\lfloor\frac{5n}{2}\right\rfloor-4$. Moreover, if $n$ is even, then the extremal planar graph is a copy of $M_{n-2}+K_{2}$; if $n$ is odd, then the extremal planar graph is either a copy of $K_{2}+M_{n-2}$, or $K_{2}\vee M_{n-2}$, or $\{u\}+O$, where $O$ represents a $C_{3}$-free outerplanar graph of even order with $ex_{\mathcal{OP}}(n-1,C_{3})=\lfloor\frac{3n}{2}\rfloor-3$ edges.
The rest of this paper is organized as follows. In Section 2, we introduce some concepts and notation on planar graphs. In Sections 3, 4 and 5, we establish sharp bounds on the planar Turán number of $H_{4},H_{5}$ and $H_{6}$, respectively, and characterize all extremal planar graphs for the latter two graphs.
@@ 2 Preliminaries
For each $H$-free TB, we can transform it into a new TB by regarding all holes whose facial cycles are $C_{3}$s with corresponding 3-faces, such that the new TB maintains $H$-free. We call such a new TB a solid TB (i.e., a solid TB contains no holes with facial cycle $C_{3}$). In the proofs of Theorems 1.1 and 1.2, we will estimate the maximum triangle-density among all $H$-free TBs and TCs (where $H$ is either $H_{4}$ or $H_{5}$). For this purpose, it suffices to estimate the maximum triangle-density among all $H$-free solid TBs and all $H$-free TCs composed of solid TBs. See Figure 2 as an example, the left side is a plane graph, whose TBs are $B_{1},B_{2}$ and $B_{3}$ (see the Figure 2 (2), two gray faces are holes). $B_{1}^{\prime}$ is a solid TB obtained from $B_{1}$ (see the Figure 2 (3)).
@@
If $\Delta_{I}(G_{E})\leq 9$, then by Ineq. (8), $e(G_{E})=\lfloor\frac{5n}{2}\rfloor-4$ if and only if $|E_{I}(G_{E})|=\frac{n}{2}+4$. By Lemma 5.3, we conclude that $e(G_{E})=\lfloor\frac{5n}{2}\rfloor-4$ implies $G_{E}$ contains a spanning subgraph $G^{\prime}$ that is a copy of $\left(\left\lfloor\frac{n-2}{2}K_{2}\right\rfloor\right)+K_{2}$. Without loss of generality, assume that $G^{\prime}=xy+M$, where $M$ is a matching of size $\lfloor(n-2)/2\rfloor$. If $n$ is even, then $e(G^{\prime})=\lfloor\frac{5n}{2}\rfloor-4=e(G)$, and hence $G=G^{\prime}$ is a copy of $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$. If $n$ is odd, then $|V(G)-V(G^{\prime})|=1$ and $|E(G)-E(G^{\prime})|=2$ (say $V(G)-V(G^{\prime})=\{z\}$ and $E(G)-E(G^{\prime})=\{e_{1},e_{2}\}$). It is clear neither $e_{1}$ nor $e_{2}$ belongs to $G[V(G^{\prime})]$, for otherwise there is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction. Hence, we can assume that $e_{1}=z_{1}z$ and $e_{2}=z_{2}z$. Then either $\{z_{1},z_{2}\}=\{x,y\}$ or $z_{1},z_{2}$ are endpoints of two edges in the matching $G^{\prime}-\{x,y\}$, respectively, for otherwise there is a $\Theta_{4}\dot{\cup}C_{3}$, a contradiction. Therefore, $G=G^{\prime}\cup\{z_{1}z,z_{2}z\}$ is either a copy of $K_{2}+\left(\frac{n-2}{2}K_{2}\right)=K_{2}+M_{n-2}$ or a copy of $K_{2}\vee M_{n-2}$. On the other hand, $\left(\frac{n-2}{2}K_{2}\right)+K_{2}$ and $K_{2}\vee M_{n-2}$ are $\Theta_{4}\dot{\cup}C_{3}$-free obviously.